if Amina called 14 times from her phone, and Deanna called 18 times from her phone, and each call took 1 ½ minutes, how many minutes total were the two girls on the phone? Give your answer in whole numbers.

Answer by tommyt3rd(5050)

(14+18)(3/2)=
32(3/2)=48

:)